OpenStudy (anonymous):

Factor completely 4x4– 20x3+ 12x2

OpenStudy (anonymous):

find x \[x ^{4}-20x ^{3}+12x ^{2}=0\]?????????????

OpenStudy (anonymous):

\[x^4-20x^3+12x^2=0 \\ x^2(x^2-20x+12)=0 \\ x^2(x^2-20x+100-100+12)=0 \\ x^2(x^2-20x+100-88)=0 \\ x^2((x-10)^2-88)=0 \\ then \ \ \ x^2=0 \ \ or \ \ (x-10)^2=88 \\ x^2=0 \ \ so \ \ x=0 \\ (x-10)^2=88 \ \ then \ \ x-10=\pm \sqrt{88} \ \ and \ \ x=10\pm \sqrt{88} \]